A population consist of the five measurements 2,6,8,3 & 1. How many different samples of size n=2 can be drawn from the population? What is the mean and variance of the sampling distribution of the sample means?
1
Expert's answer
2021-05-07T09:50:28-0400

The number of all possible sample = 5C2=10



The variance of the sampling distribution of the sample means

"s^2 = \frac{1}{10-1} \times 25.5 = 2.83"